OpenStudy (anonymous):

find indicated sum, difference, product or quotient: f(x)=16-x^2; g(x)=4-x find (f+g)(x)

OpenStudy (anonymous):

g(f(x)) = 4-(16-x^2) right so far?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so then it's 4 -16 +x^2

OpenStudy (anonymous):

and -12 +x^2

OpenStudy (anonymous):

yup! :)

OpenStudy (anonymous):

cool, now do I plug in that g(x) for f(x)= 16-x^2

OpenStudy (anonymous):

quite horrible, but just sub 4-x into the x.

OpenStudy (anonymous):

like 16 -(4-x)

OpenStudy (anonymous):

\(16 -(4-x)^2\)

OpenStudy (anonymous):

oh shoot, I forgot the exponent

OpenStudy (anonymous):

16-(4-x)^2 is 16 (-4-x)(-4-x) which is 32 +8x+x^2

OpenStudy (anonymous):

and then add that to the -12+x^2 from g(x)?

OpenStudy (anonymous):

hmm, very close, but something's off isn't it?

OpenStudy (anonymous):

16-(4-x)^2 is not 16 (-4-x)(-4-x) it's 16 - (4-x)(4-x) though.

OpenStudy (anonymous):

ohhh, that may fix it

OpenStudy (anonymous):

yeah. something will cancel out.

OpenStudy (anonymous):

yea, it left +8x -x^2

OpenStudy (anonymous):

but I still don't have it exactly right. I should be getting -x^2 +x +12 I think, but I've got a -12

OpenStudy (anonymous):

note that (f+g)(x)=16-x^2 +4-x

OpenStudy (anonymous):

(f-g)(x)=16-x^2 -(4-x)

OpenStudy (anonymous):

hmm, so what do I need to do to solve it now?

OpenStudy (anonymous):

well, first, lol which are you doing? addition or subtraction?

OpenStudy (anonymous):

well, it says f+g, so I would assume addition

OpenStudy (anonymous):

so, simplify this one and you're done. (f+g)(x)=16-x^2 +4-x

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

16-(12+x^2) +4-(12+x^2)? I think I screwed it up

OpenStudy (anonymous):

or is it 16-(+8x -x^2) +4-(+8x -x^2) I'm not too sure of anything anymore

OpenStudy (anonymous):

you don't need sub anything, just simplify as it's (f+g)(x) =f(x)+g(x)

OpenStudy (anonymous):

oh, cuz I've already done all that. duh.

OpenStudy (anonymous):

but when I add them together, I only get 20, the x^2's cancel out

OpenStudy (anonymous):

I have an option for the answer to be -x^2-x+20

OpenStudy (anonymous):

add 16-x^2 and 4-x only.

OpenStudy (anonymous):

ok, well, thats it, obviously. but what happened to the 8x and +/-2x^2's? were those just wrong?

OpenStudy (anonymous):

they weren't needed? :)

OpenStudy (anonymous):

jesus christ. haha. well ok. sorry for being such a pain in the retrice thanks for the help though

OpenStudy (anonymous):

hmm, well thats an odd substitution for @$$.... whatever lol

OpenStudy (anonymous):

lol you're welcome